OpenStudy (liizzyliizz):

find F'(x) when F(x)= integral f(t) dt ( lower limit is a upper limit is x) i know how to solve them.. but im thrown off with the whole "f(t)" help?

OpenStudy (anonymous):

f'(x)=f(x)-f(a) differential of integral gets cancelled

OpenStudy (anonymous):

GOT iT !

OpenStudy (dumbcow):

if you actually went through and integrated, you would get F = 3x - 3(0) = 3x then F' = 3

OpenStudy (anonymous):

LOL! hers f(x)=f(t)

OpenStudy (liizzyliizz):

lol he helped me on a previous one thats where he's getting that from.

OpenStudy (anonymous):

k

OpenStudy (dumbcow):

oh yeah, sorry didn't read the entire question, figured it was the same problem

OpenStudy (dumbcow):

anyway the t goes away after integrating by substituting the x in for t

OpenStudy (liizzyliizz):

yeah i figured after a while of doing several of these. It just didn't seem right. F'(x)= f(x) but ehh maybe I'm over thinking it lol

OpenStudy (dumbcow):

yeah probably, this is just basically defining what the anti-derivative is By definition, F'(x) = f(x) where F is the anti-derivative of f(x)

OpenStudy (liizzyliizz):

Yeah I understand the concept, I just like to make sure I'm not making simple mistakes.

OpenStudy (dumbcow):

:)